# How Network Topology Affects the Strength of Dangerous Power Grid Perturbations

### v2

@@ Abstract
Reasonably large perturbations may push a power grid from its stable synchronous state into an undesirable state. Identifying vulnerabilities in power grids by studying power grid stability against such perturbations can aid in preventing future blackouts. Probabilistic stability quantifiers such as basin stability, which measures the asymptotic stability of a system, and survivability, which measures the transient stability of a system, have been commonly used to quantify the stability of nodes in a power grid. However, these quantifiers do not provide information about the strength of perturbations that destabilize the system. To measure the strength of perturbations beyond which the stability of the system gets compromised, we employ two probabilistic distance-based stability measures — basin stability bound, which deals with a system’s asymptotic behaviour, and survivability bound, a newly defined stability measure that deals with a system’s transient behaviour. Using these stability quantifiers, we conduct a detailed study on the impact of network topology on the strength of dangerous power grid perturbations. We find that certain tree-like structures and node connectivity can strongly predict nodes with low asymptotic and transient stability; we highlight network structures that are particularly vulnerable.
  • September 2024
Keywords: power grids, probabilistic methods, network dynamics, stability of dynamical systems, coupled oscillator networks
@@ 1 Introduction
Power grids are critical infrastructures that underpin the functioning of modern society. Failures in these systems have resulted in large-scale blackouts that have left millions of people without electricity [1, 2, 3]. In recent times, grid expansion, modernization, and decentralization have been promoting rapid change to existing power grid infrastructure [4, 5, 6]. Thus, as power grids are becoming increasingly more complex, it is important to ensure they are resilient to various perturbations in order to prevent future blackouts.
@@
During normal operation, all parts of a power grid function at the same frequency [7]. This state is called the grid’s synchronous state. Perturbations such as a line being switched off or the power demand of the grid not being met may result in large parts of the grid desynchronizing, causing a cascading failure [8, 9, 10]. As perturbations affecting power grids can be reasonably large, linear stability analysis by means of evaluating the eigenvalues of the Jacobian matrix at an equilibrium point or the master stability function [11] cannot be employed as a measure of grid stability against such perturbations.
@@
To quantify the stability of dynamical systems such as power grids against reasonably large perturbations, several non-local stability measures have been proposed [12, 13, 14, 15, 16, 17, 18]. Among these, a popular stability measure, known as basin stability [12], relates to the fraction of phase space that forms the basin of attraction of the desirable attractor. In the context of power grids, the desirable attractor corresponds to the grid’s stable synchronous state. Basin stability has been used extensively in the study of power grid stability [19, 20, 21, 22, 23, 24, 25, 26].
@@
Basin stability deals with the asymptotic, or, long term behaviour of a system. In addition to the asymptotic behaviour, the transient behaviour, that is, the behaviour of a system before it reaches its steady state, is particularly relevant when dealing with power grids. Power grids operate within certain frequency bounds, and control mechanisms are triggered when a perturbation causes the grid to operate out of the set frequency bound. Such perturbations are undesirable for the system. A stability measure called survivability has been proposed to quantify the transient stability of dynamical systems [13]. The set of states that do not leave a given desirable region of the phase space within a given time is called the basin of survival, and survivability is the fraction of states that are part of the system’s basin of survival. Thus, survivability measures a system’s capacity to resist perturbations from amplifying past the system’s desirble region. Survivability, like basin stability, has been used to study power grid stability [13, 22, 27].
@@
Basin stability and survivability are both volume-based measures of stability as they are related to the volume of the basin of attraction and the volume of the basin of survival, respectively. For power grids, the basin of attraction of the synchronous state can be very distorted [28]. It is possible that a stronger perturbation is safe for the system, but a weaker one is not. Hence, it is crucial to understand the magnitude of perturbations that are dangerous for the system. The volume-based stability measures fail to capture this. However, basin stability bound [29], a distance-based stability measure that provides a safe bound to the strength of perturbations based on the system’s asymptotic behavior, captures this aspect. Analogous to basin stability bound, we propose a distance-based stability quantifier called survivability bound that provides a safe bound to the strength of perturbations based on the system’s transient behaviour.
@@
Using these two distance-based stability measures, we conduct a detailed study on the effect of network topology on power grid stability. Nodes with a low basin stability bound indicate that a low perturbation strength is required to cause permanent grid synchrony loss. On the other hand, nodes with a low survivability bound indicate that a low perturbation strength is required to cause undesirable transients in the system, which do not necessarily lead to loss of grid synchrony. In this study, we outline the local network topology that results in low single-node stability.
@@ 2.1 Power grid model
We use a complex network representation to model power grids with generators and consumers as nodes and transmission lines as edges. Generators and consumers are modelled as synchronous machines that follow the swing equation [7]. The equations that describe the dynamics of the grid are [30]
$\dot{\theta_{i}}$ $=$ $\omega_{i},$ (1a)
$I_{i}\dot{\omega_{i}}$ $=$ $-\gamma_{i}\omega_{i}+\bar{P_{i}}-\sum^{N}_{j=1}T_{ij}\sin(\theta_{i}-\theta_{j}),$ (1b)
@@
$\dot{\theta_{i}}$ $=$ $\omega_{i},$ (1ba)
$\dot{\omega_{i}}$ $=$ $-\alpha\omega_{i}+P_{i}-\sum^{N}_{j=1}K_{ij}\sin(\theta_{i}-\theta_{j}),,$ (1bb)
@@
In this paper, we use the simplified differential equations described by equations (1ba) and (1bb) to model power grids. Each node in the network has two corresponding dynamical variables — a phase angle and a frequency. The fixed point of these equations corresponds to the stable synchronous state of the grid. In this state, the $i^{\rm th}$ node has a phase $\theta_{i}^{s}$ and frequency 00 in the rotating frame of reference.
@@ 2.2 Basin stability bound
We first describe the basin stability bound introduced in [12, 19].
@@
$\beta(X^{P})=\frac{\textnormal{Vol}(\mathcal{B}\cap X^{P})}{\textnormal{Vol}(X^{P})}$ (1bc)
@@
We consider a finite subset of the phase space $X^{0}\subseteq$ X, representing the extent to which perturbations can push the system. Basin stability bound is defined in this region $X^{0}$.
@@
${X^{D}}(d)=\{x\in{X^{0}}\>|\>\textnormal{dist}(x,\mathcal{A})<d\}$ (1bd)
@@
${D}$ is the set of distances at which the corresponding basin stability is less than a basin stability tolerance $t_{\beta}$. Thus,
${D_{\beta}}=\{d\in(0,d_{\textnormal{max}}]\>|\>\beta({X^{D}}(d))<t_{\beta}\}$ (1be)
@@
The basin stability bound of the attractor $\mathcal{A}$ is defined as [29]
$\bar{\beta}=\cases{\textnormal{inf}(D_{\beta})&if $D_{\beta}\neq\phi$\\ d_{\textnormal{max}}&otherwise\\ }$ (1bf)
@@
$X_{i}^{0}=\{(\theta,\omega)\in X|(\theta_{i}\in X^{sn}\land\omega_{i}=0\land(\forall j\neq i:\theta_{j}=\theta^{s}_{j}\land\omega_{j}=0)\}$ (1bg)
@@
$\hat{\beta}(X^{P})=\frac{n_{\beta}}{n}$ (1bh)
To compute the basin stability bound, the basin stability $\hat{\beta}(X^{D}(d))$ is computed from $d=d_{\textnormal{max}}$ to $d=d_{0}$, using $n$ samples for every basin stability estimation. $d_{0}$ is the largest distance such that $\hat{\beta}({X^{D}}(d_{0}))=1$. For $\hat{\beta}(X^{D}(d))<t_{\beta}$, the values of $d$ are noted and added to a set ${D_{\beta}}$. The basin stability bound is computed using equation 1bf. Refer to the paper by Alvares et al. [29] for a detailed computation procedure and the error associated in basin stability bound’s estimation.
@@ 2.3 Survivability bound
Suppose the system has a desirable region $X^{+}\subseteq X$ of the phase space. A perturbation is considered safe if the state does not leave the desirable region in a finite time $t$. Let $X^{S}_{t}$ be the set of points in $X^{P}$ that do not leave the desirable region $X^{+}$ in the time $t$. Assuming a uniform distribution of perturbations in the region $X^{P}$, survivability is defined as [13, 22]
$\sigma(X^{P})=\frac{\textnormal{Vol}(X_{t}^{S})}{\textnormal{Vol}(X^{0})}$ (1bi)
@@
${D_{\sigma}}=\{d\in(0,d_{\textnormal{max}}]\>|\>\sigma({X^{D}}(d))<t_{\sigma}\}$ (1bj)
where $t_{\sigma}\in(0,1]$ is a predefined survivability tolerance, $d_{\textnormal{max}}$ is the maximum distance we would like to consider, and $X^{D}(d)$ is given by equation (1bd).
@@
$\bar{\sigma}=\cases{\textnormal{inf}(D_{\sigma})&if $D_{\sigma}\neq\phi$\\ d_{\textnormal{max}}&otherwise\\ }$ (1bk)
@@
In the case of the power grids, the desirable region is defined to be
$X^{+}=\{(\theta,\omega)\in X|\forall i:-\pi<\theta_{i}<\pi\land-\omega^{+}<\omega_{i}<\omega^{+}\}$ (1bl)
@@
If single-node perturbations occur in the region $\theta_{i}\in X^{sn}$, then the single-node survivability bound of node $i$ is the survivability bound defined in the phase space region given by equation (1bg).
Survivability is estimated using a Monte Carlo simulation. In the region ${X^{P}}$, a number of initial conditions, $n$, are uniformly sampled. If the number of initial conditions that converge to the attractor $\mathcal{A}$ is $n_{\sigma}$, then the estimated survivability is
$\hat{\sigma}=\frac{n_{\sigma}}{n}$ (1bm)
The survivability bound is computed using the same procedure used to compute the basin stability bound.
@@ 2.4 Quantifying the strength of power grid perturbations
Basin stability bound and survivability bound rely on a notion of distance between a perturbed state and the attractor. This distance is indicative of the strength of perturbation. Distance in the phase space of a dynamical system can be quantified in various ways, ranging from standard Euclidean distance to the energy difference between the two states [29, 31]. For the distance of a perturbation from the synchronous state to have a physical meaning, we quantify it by defining an energy function that is related to the energy change of the system due to the perturbation.
Note that we have used a simplified power grid model given by equations (1ba) and (1bb) with the inertia at all nodes being the same and quantities such as torque and energy that we refer to are normalized by inertia.
@@
[…]
@@
$\tau(\theta_{i})=\cases{-\textnormal{min}\{0,\frac{d^{2}\theta_{i}}{dt^{2}}\}&if $\theta_{i}^{p}>\theta_{i}>\theta_{i}^{s}$\\ -\textnormal{max}\{0,\frac{d^{2}\theta_{i}}{dt^{2}}\}&if $\theta_{i}^{p}<\theta_{i}<\theta_{i}^{s}$}$ (1bo)
Additionally, we also assume that the perturbation is made at the synchronous grid frequency, which makes $w_{i}=0$. The torque required for this is
$\tau(\theta_{i})=\cases{-\textnormal{min}\{0,(P_{i}-\sum^{N}_{j=1}K_{ij}\sin(\theta_{i}-\theta_{j}))\}&if $\theta_{i}^{p}>\theta_{i}>\theta_{i}^{s}$\\ -\textnormal{max}\{0,(P_{i}-\sum^{N}_{j=1}K_{ij}\sin(\theta_{i}-\theta_{j}))\}&if $\theta_{i}^{p}<\theta_{i}<\theta_{i}^{s}$}$ (1bp)
@@
$E(\theta_{i}^{p})=\cases{\textnormal{min}\{W(\theta_{i}^{s},\theta_{i}^{p}),\>W(\theta_{i}^{s},-2\pi+\theta_{i}^{p})\}&if $\theta_{i}^{p}>\theta_{i}^{s}$\\ \textnormal{min}\{W(\theta_{i}^{s},\theta_{i}^{p}),\>W(\theta_{i}^{s},2\pi+\theta_{i}^{p})\}&if $\theta_{i}^{p}<\theta_{i}^{s}$}$ (1bq)
@@
$\textnormal{dist}(x_{i}^{p},\mathcal{A})=E(\theta_{i}^{p})$ (1br)
[…]

### v3

@@ Abstract
Reasonably large perturbations may push a power grid from its stable synchronous state into an undesirable state. Identifying vulnerabilities in power grids by studying power grid stability against such perturbations can aid in preventing future blackouts. Probabilistic stability quantifiers such as basin stability, which measures the asymptotic stability of a system, and survivability, which measures the transient stability of a system, have been commonly used to quantify the stability of nodes in a power grid. However, these quantifiers do not provide information about the strength of perturbations that destabilize the system. To measure the strength of perturbations beyond which the stability of the system gets compromised, we employ two probabilistic distance-based stability measures — basin stability bound, which deals with a system’s asymptotic behaviour, and survivability bound, a newly defined stability measure that deals with a system’s transient behaviour. Using these stability quantifiers, we conduct a detailed study on the impact of network topology on the strength of dangerous power grid perturbations. In this, we uncover a new class of highly vulnerable nodes that were previously unknown. Additionally, we establish connections with tree-like network structures and node connectivity to lowly stable nodes.
keywords:
Network stability, basin stability, nonlinear oscillations, power grids, Kuramoto model
@@ 1 Introduction
Power grids are critical infrastructures that underpin the functioning of modern society. Failures in these systems have resulted in large-scale blackouts that have left millions of people without electricity Central Electricity Regulatory Commission , 2012 [CERC]; Project Group Turkey, [2015]; Union for the Coordination of Transmission of Electricity , 2003 [UCTE]. In recent times, expansion, modernization, and decentralization of the grid have been promoting rapid changes in existing power grid infrastructure Pepermans et al., [2005]; Di Silvestre et al., [2018]; Fang et al., [2011]. Thus, as power grids become increasingly complex, it is important to ensure that they are resilient to various perturbations to prevent future blackouts.
@@
During normal operation, all parts of a power grid function at the same frequency Machowski et al., [2020]. This state is called the grid’s synchronous state. Perturbations such as a line being switched off or the power demand of the grid not being met may result in large parts of the grid desynchronizing, causing cascading failure Buldyrev et al., [2010]; Motter and Lai, [2002]; Schäfer et al., [2018]. As disturbances affecting power grids can be reasonably large, linear stability analysis by means of evaluating the eigenvalues of the Jacobian matrix at an equilibrium point or the master stability function Pecora and Carroll, [1998] cannot be employed as a measure of grid stability against such perturbations.
@@
To quantify the stability of dynamical systems, such as power grids, against reasonably large perturbations, several non-local stability measures have been proposed Menck et al., [2013]; Hellmann et al., [2016]; Mitra et al., [2015, 2017]; Klinshov et al., [2015]; Halekotte and Feudel, [2020]; Klinshov et al., [2018]. Among these, a popular stability measure, known as basin stability Menck et al., [2013], relates to the fraction of phase space that forms the basin of attraction of the desirable attractor. In the context of power grids, the desirable attractor corresponds to the stable synchronous state of the grid. Basin stability has been used extensively in the study of power grid stability Menck et al., [2014]; Ji and Kurths, [2014]; Schultz et al., 2014a ; Nitzbon et al., [2017]; Kim et al., [2016]; Nauck et al., [2022]; Feld and Hartmann, [2019]; Schäfer et al., [2016].
@@
Basin stability deals with the asymptotic, or, long-term behavior of a system. In addition to asymptotic behavior, transient behavior, that is, the behavior of a system before it reaches its steady state, is particularly relevant when dealing with power grids. Power grids operate within certain frequency bounds, and control mechanisms are triggered when a disturbance causes the grid to operate outside the set frequency bound. Such perturbations are undesirable for the system. A stability measure called survivability has been proposed to quantify the transient stability of dynamical systems Hellmann et al., [2016]. The set of states that do not leave a given desirable region of the phase space within a given time is called the basin of survival, and survivability is the fraction of states that are part of the system’s basin of survival. Thus, survivability measures a system’s capacity to resist perturbations from amplifying past the system’s desirble region. Survivability, like basin stability, has been used to study the stability of the electricity grid Hellmann et al., [2016]; Nitzbon et al., [2017]; Büttner et al., [2022].
@@
Basin stability and survivability are both volume-based measures of stability as they are related to the volume of the basin of attraction and the volume of the basin of survival, respectively. For power grids, the basin of attraction of the synchronous state can be very distorted Halekotte et al., [2021]. It is possible that a stronger perturbation is safe for the system, but a weaker one is not. Hence, it is crucial to understand the magnitude of perturbations that are dangerous to the system. Volume-based stability measures fail to capture this. However, basin stability bound Alvares and Banerjee, [2024], a distance-based stability measure that provides a safe bound to the strength of perturbations based on the asymptotic behavior of the system, captures this aspect. Analogous to basin stability bound, we propose a distance-based stability quantifier called survivability bound that provides a safe bound to the strength of perturbations based on the system’s transient behavior. Although, in this paper, we use the above methods on power grids, they are readily applicable to other dynamical systems.
@@
Using the two distance-based stability measures, we conducted a detailed study on the effect of network topology on the stability of the power grid. Nodes with a low basin stability bound indicate that a low perturbation strength is required to cause permanent loss of grid synchrony. On the other hand, nodes with a low survivability bound indicate that a low perturbation strength is required to cause undesirable transients in the system, which do not necessarily lead to loss of grid synchrony. In this study, based on distance-based stability quantifiers and a novel method of quantifying the strength of power grid perturbations, we outline the local network topology that results in low single-node stability and uncover a previously unknown, new class of highly vulnerable nodes.
@@ 2.1 Power grid model
We use a complex network representation to model power grids with generators and consumers as nodes and transmission lines as edges. Generators and consumers are modeled as synchronous machines that follow the swing equation Machowski et al., [2020]. The equations that describe the dynamics of the grid are Filatrella et al., [2008]
$\dot{\theta_{i}}$ $=$ $\omega_{i},$ (1)
$I_{i}\dot{\omega_{i}}$ $=$ $-\gamma_{i}\omega_{i}+\bar{P_{i}}-\sum^{N}_{j=1}T_{ij}\sin(\theta_{i}-\theta_{j}),$ (2)
@@
$\dot{\theta_{i}}$ $=$ $\omega_{i},$ (3)
$\dot{\omega_{i}}$ $=$ $-\alpha\omega_{i}+P_{i}-\sum^{N}_{j=1}K_{ij}\sin(\theta_{i}-\theta_{j}),,$ (4)
@@
In this paper, we use the simplified differential equations described by equations (3) and (4) to model power grids. Each node in the network has two corresponding dynamical variables — a phase angle and a frequency. The fixed point of these equations corresponds to the stable synchronous state of the grid. In this state, the $i^{\rm th}$ node has a phase $\theta_{i}^{s}$ and frequency 0 in the rotating frame of reference.
@@ 2.2 Basin stability bound
We first describe the basin stability introduced in Menck et al., [2013, 2014].
@@
$\beta(X^{P})=\frac{\textnormal{Vol}(\mathcal{B}\cap X^{P})}{\textnormal{Vol}(X^{P})}$ (5)
@@
To define basin stability bound Alvares and Banerjee, [2024], we consider a finite subset of the phase space $X^{0}\subseteq$ X, representing the extent to which perturbations can push the system. Basin stability bound is defined in this region $X^{0}$.
@@
${X^{D}}(d)=\{x\in{X^{0}}\>|\>\textnormal{dist}(x,\mathcal{A})<d\}$ (6)
@@
$D_{\beta}$} is the set of distances at which the corresponding basin stability is less than a basin stability tolerance $t_{\beta}$. Thus,
${D_{\beta}}=\{d\in(0,d_{\textnormal{max}}]\>|\>\beta({X^{D}}(d))<t_{\beta}\}$ (7)
@@
The basin stability bound of the attractor $\mathcal{A}$ is defined as Alvares and Banerjee, [2024]
$\bar{\beta}=\begin{cases}\inf(D_{\beta})&\text{if }D_{\beta}\neq\varnothing\\ d_{\text{max}}&\text{otherwise}\end{cases}$ (8)
@@
$X_{i}^{0}=\{(\theta,\omega)\in X|(\theta_{i}\in X^{sn}\land\omega_{i}=0\land$ (9)
$(\forall j\neq i:\theta_{j}=\theta^{s}_{j}\land\omega_{j}=0)\}$ (10)
@@
$\hat{\beta}(X^{P})=\frac{n_{\beta}}{n}$ (11)
l
To compute the basin stability bound, the basin stability $\hat{\beta}(X^{D}(d))$ is computed from $d=d_{\textnormal{max}}$ to $d=d_{0}$, using $n$ samples for every basin stability estimation. $d_{0}$ is the largest distance such that $\hat{\beta}({X^{D}}(d_{0}))=1$. For $\hat{\beta}(X^{D}(d))<t_{\beta}$, the values of $d$ are noted and added to a set $D_{\beta}$}. The basin stability bound is computed using equation 8. Refer to the paper by Alvares et al. Alvares and Banerjee, [2024] for a detailed computation procedure and the error associated in basin stability bound’s estimation.
@@ 2.3 Survivability bound
Suppose that the system has a desirable region $X^{+}\subseteq X$ of the phase space. A perturbation is considered safe if the state does not leave the desirable region in finite time $t$. Let $X^{S}_{t}$ be the set of points in $X^{P}$ that do not leave the desirable region $X^{+}$ in the time $t$. Assuming a uniform distribution of perturbations in the region $X^{P}$, survivability is defined as Hellmann et al., [2016]; Nitzbon et al., [2017]
$\sigma(X^{P})=\frac{\textnormal{Vol}(X_{t}^{S})}{\textnormal{Vol}(X^{0})}$ (12)
@@
${D_{\sigma}}=\{d\in(0,d_{\textnormal{max}}]\>|\>\sigma({X^{D}}(d))<t_{\sigma}\}$ (13)
where $t_{\sigma}\in(0,1]$ is a predefined survivability tolerance, $d_{\textnormal{max}}$ is the maximum distance we would like to consider, and $X^{D}(d)$ is given by equation (6).
@@
$\bar{\sigma}=\begin{cases}\inf(D_{\sigma})&\text{if }D_{\sigma}\neq\varnothing\\ d_{\text{max}}&\text{otherwise}\end{cases}$ (14)
@@
In the case of power grids, the desirable region is defined to be
$X^{+}=\{(\theta,\omega)\in X|\forall i:-\pi<\theta_{i}<\pi\land-\omega^{+}<\omega_{i}<\omega^{+}\}$ (15)
@@
If single-node perturbations occur in the region $\theta_{i}\in X^{sn}$, then the single-node survivability bound of node $i$ is the survivability bound defined in the phase space region given by equation (9).
Survivability is estimated using a Monte Carlo simulation. In the region $X^{P}$}, a number of initial conditions, $n$, are uniformly sampled. If the number of initial conditions that converge to attractor $\mathcal{A}$ is $n_{\sigma}$, then the estimated survivability is
$\hat{\sigma}=\frac{n_{\sigma}}{n}$ (16)
The survivability bound is computed using the same procedure as used to compute the basin stability bound.
@@ 2.4 Quantifying the strength of power grid perturbations
Basin stability bound and the survivability bound rely on the notion of distance between a perturbed state and the attractor. This distance is indicative of the strength of the perturbation. The distance in the phase space of a dynamical system can be quantified in various ways, ranging from standard Euclidean distance to the energy difference between the two states Alvares and Banerjee, [2024]; Lundström, [2018]. For the distance of a perturbation from the synchronous state to have a physical meaning, we quantify it by defining an energy function that is related to the energy change of the system due to the perturbation.
Note that we have used a simplified power grid model given by equations (3) and (4) with the inertia at all nodes being the same and quantities such as torque and energy that we refer to are normalized by inertia.
@@
$W(\theta_{i}^{s},\theta_{i}^{p})=\int^{\theta_{i}^{p}}_{\theta_{i}^{s}}\tau(\theta_{i})\>d\theta_{i}$ (17)
@@
$\tau(\theta_{i})=\begin{cases}-\textnormal{min}\{0,\frac{d^{2}\theta_{i}}{dt^{2}}\}&\textnormal{if }\theta_{i}^{p}>\theta_{i}>\theta_{i}^{s}\\ -\text{max}\{0,\frac{d^{2}\theta_{i}}{dt^{2}}\}&\text{if }\theta_{i}^{p}<\theta_{i}<\theta_{i}^{s}\end{cases}$ (18)
Additionally, we also assume that the perturbation is made at the synchronous grid frequency, which makes $w_{i}=0$. With this assumption,
$\frac{d^{2}\theta_{i}}{dt^{2}}=P_{i}-\sum^{N}_{j=1}K_{ij}\sin(\theta_{i}-\theta_{j})$ (19)
@@
$E(\theta_{i}^{p})=\begin{cases}\textnormal{min}\{W(\theta_{i}^{s},\theta_{i}^{p}),\>W(\theta_{i}^{s},-2\pi+\theta_{i}^{p})\}&\text{if }\theta_{i}^{p}>\theta_{i}^{s}\\ \text{min}\{W(\theta_{i}^{s},\theta_{i}^{p}),\>W(\theta_{i}^{s},2\pi+\theta_{i}^{p})\}&\text{if }\theta_{i}^{p}<\theta_{i}^{s}\end{cases}$ (20)
@@
$\textnormal{dist}(x_{i}^{p},\mathcal{A})=E(\theta_{i}^{p})$ (21)
where $x_{i}^{p}=(\theta_{1}^{s},\>\theta_{2}^{s},\>...,\>\theta_{i}^{p},\>...,\>\theta_{N}^{s},\>0,\>0,\>...,\>0,\>...,\>0)$